# How AI Coding Agents Modify Code: A Large-Scale Study of GitHub Pull Requests

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: MSR ’26: Proceedings of the 23rd International Conference on Mining Software Repositories; April 2026; Rio de Janeiro, Brazilccs: Software and its engineering Collaboration in software developmentccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Artificial intelligence (AI) coding agents are rapidly transforming software development. Tools such as GitHub Copilot (GitHub Copilot, 2025), OpenAI Codex (OpenAI, 2025), Claude Code (Anthropic, 2025), Cursor (Cursor, 2025), and Devin (Devin AI, 2025) can now autonomously generate code, fix bugs, and submit pull requests (PRs). These systems are evolving from assistive tools into active collaborators, a transition towards Software Engineering 3.0 (Hassan et al., 2025, 2024). Prior work has examined how developers interact with AI-generated suggestions and how such tools affect productivity and code quality (Chen et al., 2021; Vaithilingam et al., 2022; Li et al., 2025; Barke et al., 2023; Hasan et al., 2025; Vaithilingam et al., 2023; Ogenrwot and Businge, 2025), highlighting both opportunities and challenges for human–AI collaboration.
@@ 2. Method
Step 3: PRs Filtering Criteria. We focus on merged PRs, which provide complete commit histories. For Agentic PRs, we removed entries lacking patch text, required for computing structural and similarity metrics. A commit was considered to have a valid patch if the GitHub API returned a non-empty unified diff for all modified files. For Human PRs, we excluded cases with missing repositories, unavailable commit data, or incomplete patches retrieved from the GitHub API. After filtering, the analysis set contains 24,014 Agentic PRs (440,295 commits with valid patches) and 5,081 Human PRs (23,242 commits with valid patches). These datasets form the basis of all analyses in RQ1 and RQ2. Agentic PRs are also grouped by agent type for stratified analyses in the results.
@@ 3. Results
We first examine the distribution of lexical and semantic similarity scores for Human and Agentic PRs. Figure 4 shows a clear separation between lexical and semantic similarity. Lexical scores for both Agentic and Human PRs cluster near zero, indicating very limited surface-level vocabulary overlap between descriptions and diffs. In contrast, semantic similarity scores form tight peaks between 0.9 and 1.0, suggesting that both description types capture the underlying meaning of their patches even when exact wording differs.
Figure 5 presents distributional comparisons across all four metrics. Agentic PRs show a slight right-shift relative to Human PRs across both lexical and semantic measures, although the difference is small in absolute terms and both groups show high semantic alignment. The lexical measures remain heavily skewed, while the semantic models produce consistently high similarity scores, with Agentic PRs displaying tighter clustering. Okapi BM25 exhibits extremely high variance and large positive or negative values, particularly for Agentic PRs, reflecting the unbounded nature of the score and its sensitivity to tokenization. As a result, Okapi BM25 should be interpreted only as a coarse lexical indicator rather than a calibrated similarity measure.
@@ 5. Related Work
Research on AI-assisted programming has examined models such as Codex and GitHub Copilot, focusing on code completion, developer productivity, and interaction patterns (Chen et al., 2021; Vaithilingam et al., 2022, 2023; Barke et al., 2023; Ogenrwot and Businge, 2024, 2025). Most of this work investigates snippet-level behavior rather than project-scale contributions such as pull requests.
Software engineering studies on human-authored patches have analyzed patch size, structure, and review effort (Gousios et al., 2014; Kononenko et al., 2016; Wang et al., 2025; Watanabe et al., 2025) and their relationships to maintainability and defect risk (Rahman et al., 2013; Chen and Jiang, 2025). Empirical analyses of AI-generated patches remain limited, though early work reports refactoring (Horikawa et al., 2025), stylistic repetition, boilerplate generation, and occasional inconsistencies between descriptions and code changes (Tambon et al., 2025; Horikawa et al., 2025; Wu et al., 2025). The AIDev dataset (Li et al., 2025) enables the first large-scale, PR-level analysis of how AI agents author and describe code changes in real-world repositories, addressing a gap in existing literature.

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2026conference: 23rd International Conference on Mining Software Repositories; April 13–14, 2026; Rio de Janeiro, Brazilbooktitle: 23rd International Conference on Mining Software Repositories (MSR ’26), April 13–14, 2026, Rio de Janeiro, Brazilccs: Software and its engineering Collaboration in software developmentccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Artificial intelligence (AI) coding agents are rapidly transforming software development. Tools such as GitHub Copilot (GitHub Copilot, 2025), OpenAI Codex (OpenAI, 2025), Claude Code (Anthropic, 2025), Cursor (Cursor, 2025), and Devin (Devin AI, 2025) can now autonomously generate code, fix bugs, and submit pull requests (PRs). These systems are evolving from assistive tools into active collaborators, a transition towards Software Engineering 3.0 (Hassan et al., 2025, 2024). Prior work has examined how developers interact with AI-generated suggestions and how such tools affect productivity and code quality (Vaithilingam et al., 2022; Li et al., 2025; Vaithilingam et al., 2023; Ogenrwot and Businge, 2025a), highlighting both opportunities and challenges for human–AI collaboration.
@@ 2. Method
Step 3: PR Filtering. We focus on merged PRs, which provide complete commit histories. For Agentic PRs, we removed entries lacking patch text, required for computing structural and similarity metrics. A commit was considered to have a valid patch if the GitHub API returned a non-empty unified diff for all modified files. For Human PRs, we excluded cases with missing repositories, unavailable commit data, or incomplete patches retrieved from the GitHub API. After filtering, the analysis set contains 24,014 Agentic PRs (440,295 commits with valid patches) and 5,081 Human PRs (23,242 commits with valid patches). These datasets form the basis of all analyses in RQ1 and RQ2. Agentic PRs are also grouped by agent type for stratified analyses in the results, following established large-scale mining practices (Businge et al., 2022b, a, 2020, 2018).
@@ 3. Results
We first examine the distribution of lexical and semantic similarity scores for Human and Agentic PRs. Figure 4 shows a clear separation between lexical and semantic similarity. Lexical scores for both Agentic and Human PRs cluster near zero, indicating very limited surface-level vocabulary overlap between descriptions and diffs. In contrast, semantic similarity scores form tight peaks between 0.9 and 1.0, suggesting that both description types capture the underlying meaning of their patches even when wording differs.
Figure 5 presents distributional comparisons across all four metrics. Agentic PRs show a slight right shift relative to Human PRs across both lexical and semantic measures, although differences are small and both groups exhibit high semantic alignment. Lexical measures remain heavily skewed, while semantic models produce consistently high similarity scores, with Agentic PRs showing tighter clustering. Okapi BM25 exhibits very high variance and large positive or negative values, particularly for Agentic PRs, reflecting its unbounded nature and sensitivity to tokenization. As a result, Okapi BM25 should be interpreted only as a coarse lexical indicator rather than a calibrated similarity measure.
@@ 5. Related Work
Research on AI-assisted programming has examined models such as Codex and GitHub Copilot, focusing on code completion, developer productivity, and interaction patterns (Vaithilingam et al., 2022, 2023; Barke et al., 2023; Ogenrwot and Businge, 2025a, 2024). Most of this work investigates snippet-level behavior rather than project-scale contributions such as pull requests.
Software engineering studies on human-authored patches have analyzed patch size, structure, and review effort (Gousios et al., 2014; Kononenko et al., 2016; Wang et al., 2025; Watanabe et al., 2025; Ogenrwot and Businge, 2025b) and their relationships to maintainability and defect risk (Rahman et al., 2013; Chen and Jiang, 2025). Empirical analyses of AI-generated patches remain limited, though early work reports refactoring (Horikawa et al., 2025), stylistic repetition, boilerplate generation, and occasional inconsistencies between descriptions and code changes (Tambon et al., 2025; Horikawa et al., 2025; Wu et al., 2025). The AIDev dataset (Li et al., 2025) enables the first large-scale, PR-level analysis of how AI agents author and describe code changes in real-world repositories, addressing a gap in existing literature.